# AVA-Bench: Atomic Visual Ability Benchmark for Vision Foundation Models

### v2

@@ 1 Introduction
Figure 2: Visual Question Answering (VQA) often requires multiple atomic visual abilities (AVAs) to answer a question. As such, when a model makes an incorrect prediction, it can be difficult to determine whether it stems from a failure to capture all required AVAs or just a single critical one.
Vision Foundation Models (VFMs), pre-trained on large and diverse datasets, have become central to AI by providing transferable features for a wide range of downstream tasks khan2022transformers ; awais2025foundation ; chowdhury2025prompt ; bommasani2021opportunities . The variety of pre-training objectives and supervision signals has led to a proliferation of specialized VFMs—such as DINOv2 oquab2023dinov2 , CLIP radford2021learning , and SAM kirillov2023segment —each excelling in distinct visual capabilities while often exhibiting interesting emergent properties goldblum2023battle ; caron2021emerging ; naseer2021intriguing . Consequently, establishing a systematic and effective evaluation protocol for VFMs has become increasingly crucial.
Existing evaluation protocols can generally be categorized into two groups. The first focuses on task-specific capabilities, typically attaching tailored heads to VFMs, followed by fine-tuning and evaluation on dedicated datasets such as ImageNet for classification han2022survey ; mai2024fine and COCO for detection or segmentation thisanke2023semantic ; balachandran2024eureka . To better capture the diverse and complex perception challenges of the real world, recent studies advocate a more generic approach that leverages large language models (LLMs) as general-purpose heads, evaluating VFMs on broad Visual Question Answering (VQA) benchmarks liu2023visual ; zhu2023minigpt ; chowdhery2023palm .
While increasingly adopted, this generic protocol suffers from two potential blind spots: (i) discrepancies between the instruction tuning data and test VQA data can cause performance drops due to data mismatch rather than genuine visual limitations in VFMs; and (ii) existing VQA benchmarks typically require multiple visual abilities simultaneously, making it difficult to determine whether a failure arises from the absence of multiple abilities or merely a single critical one.
To address these challenges, we introduce AVA-Bench, the first VFM evaluation benchmark explicitly designed to disentangle Atomic Visual Abilities (AVAs)—the fundamental visual capabilities that combine to solve complex visual reasoning tasks. For instance, answering typical VQA questions shown in Figure 2 necessitates integrating several AVAs.
@@
We systematically benchmark leading VFMs trained under diverse objectives and data (see Figure 1), covering language-supervised (e.g., SigLIP-1/2 tschannen2025siglip ; zhai2023sigmoid , CLIP radford2021learning , InternVL-2.5 chen2024expanding ), multimodal autoregressive (AIMv2 fini2024multimodal ), segmentation-supervised (SAM kirillov2023segment ), depth-supervised (MiDas ranftl2020towards ), contrastive self-supervised (DINOv2 oquab2023dinov2 ), and agglomerative models (RADIO ranzinger2024radio ). We follow the standard protocol of adding an LLM on top, but fine-tune it separately for each AVA.
@@
  • We release a resource-efficient evaluation protocol along with an open-source codebase to facilitate the development of the next generation of accountable and versatile VFMs.
2 Vision Foundation Models (VFMs) and Evaluation
Figure 3: AVA-Bench consists of 14 Atomic Visual Abilities (AVAs) that can be combined to address more complex visual reasoning tasks.
@@ 2.1 Vision Foundation Models (VFMs)
A wide range of architectures and learning strategies have been explored to develop vision foundation models (VFMs). In general, they fall into two categories based on the nature of their training data. On one hand, VFMs are pre-trained purely on visual data. For instance, the Vision Transformer (ViT)  dosovitskiy2020image is trained on labeled images using the supervised signal to capture visual representations. DINOv2  oquab2023dinov2 , in contrast, adopts a self-supervised learning approach, enabling the model to learn those features without labeling. Some VFMs are designed for specific vision tasks—for example, SAM kirillov2023segment specializes in open-vocabulary segmentation, while MiDaS ranftl2020towards focuses on monocular depth estimation. RADIO  ranzinger2024radio instead introduces a multi-teacher distillation framework that unifies the strengths of different VFMs (e.g., CLIP, DINOv2, and SAM) into a single efficient student model.
Another family of VFMs leverages image-text pairs. CLIP  radford2021learning uses contrastive learning to align image and textual descriptions, while SigLIP  tschannen2025siglip ; zhai2023sigmoid replaces the contrastive loss with a sigmoid one for more efficient training. Unlike the conventional language-guided VFMs that process images and text separately, AIMv2  fini2024multimodal integrates visual and textual understanding into a single auto-regressive framework, which is simple yet effective. Due to their different architectures and learning objectives, VFMs may have varying strengths and limitations in visual understanding.
@@ 2.2 LLM-based Evaluation for VFMs
Figure 4: Evaluation pipeline for AVA-Bench: The standard LLaVA-style two-stage training prepares the connector and LLM for VFM evaluation. For each AVA, only connector and LoRA is trained.
Figure 5: Overall statistics of AVA-Bench.
Unlike the traditional paradigm, where different perception tasks, such as classification, detection or segmentation, necessitate task-specific models han2022survey ; awais2025foundation , the widespread adoption of Large Language Models (LLMs) as versatile interfaces has significantly shifted this paradigm zhang2024mm ; wu2023multimodal . Reflecting this shift, recent studies advocate leveraging large language models (LLMs) as general-purpose heads and evaluating VFMs on broad Visual Question Answering (VQA) benchmarks tong2024cambrian . Specifically, following the LLaVA approach, these studies utilize a two-stage framework: (i) pre-training a connector between frozen LLM and VFMs with image-text pairs for feature alignment (ii) fine-tuning both the connector and the LLM using instruction-tuning data, while keeping VFMs frozen during both stages (see Figure 5). This LLM-based evaluation protocol has rapidly gained popularity as it closely mirrors the contemporary multimodal LLM setting and effectively captures diverse real-world perception challenges.
However, this evaluation protocol suffers from two potential blind spots. First, discrepancies may exist between instruction-tuning datasets and the test VQA datasets. Consequently, a miss-prediction might arise from data mismatch rather than genuine visual deficiencies in VFMs. Second, typical VQA benchmarks typically require multiple visual abilities simultaneously to produce correct answers. This multi-ability requirement obscures the exact cause behind a model’s incorrect predictions. For instance, if a VFM is proficient in almost all visual abilities except orientation recognition (forward or backward), it would fail on the question in Figure 2 (left). In such scenarios, current evaluation methods only provide relative comparisons between VFMs without elucidating the specific missing visual capabilities. Thus, our AVA-Bench aims to complement the popular LLM-based evaluation by pinpointing exactly where a VFM excels or falters, yielding a holistic understanding of their strengths and shortcomings.
@@ 3 AVA-Bench
To address the blind spots discussed earlier, we introduce AVA-Bench, the first systematic evaluation suite explicitly designed to disentangle 14 fundamental perceptual skills–Atomic Visual Abilities (AVAs)–for VFMs. We first introduce the concept of AVAs (subsection 3.1), followed by detailing our dataset curation pipeline (subsection 3.2). Lastly, we outline the overall dataset statistics and discuss quality control (subsection 3.3).
@@ 3.1 Atomic Visual Abilities (AVAs)
AVAs are fundamental perceptual capabilities that can be combined to address more complex visual reasoning tasks. Rather than treating VQA as a monolithic task, we break it down into 14 fundamental perceptual abilities–such as counting, depth estimation, localization and spatial reasoning–each of which can be composed to answer more complex questions. For example, answering the question, “How many yellow dogs are facing backward on the left-hand side of the stop sign?”, requires multiple AVAs: counting, color recognition, localization, and spatial reasoning (see Figure 2). By explicitly defining the set of AVAs required to interpret a question, our benchmark quantitatively characterizes where a VFM excels or falters across each core ability.
The AVAs selected for AVA-Bench are grounded in a thorough literature analysis, including compositional text-to-images benchmarks huang2023t2i ; kil2024mllm ; wu2024conceptmix and VQA questions goyal2017making ; RealWorldQA ; ainslie2023gqa (details in Appendix), focusing strictly on pure perceptual tasks and excluding non-perceptual reasoning skills (e.g., historical context, mathematical reasoning). Examples of each AVA can be found in Figure 3 and detailed definitions can be found in Appendix.
@@ 3.2 Dataset Curation
Constructing AVA-Bench required carefully isolating image–question pairs that specifically test individual AVAs. Existing MLLM and VQA datasets often blend multiple perceptual abilities, complicating direct assessments tong2024eyes ; zhang2024mme ; yu2023mm . To achieve clear isolation, we assembled a comprehensive suite comprising image–question pairs from 26 diverse datasets, explicitly aligning each pair with a single targeted AVA. These datasets span a broad range of domains—including general scenes, wildlife (e.g., birds, fungi, plants), vehicles, indoor/outdoor settings, and remote-sensing imagery.
To construct the benchmark for an AVA, image–question pairs were carefully designed or adapted to focus solely on that AVA. For instance, the original question “What’s the depth of the car from the camera?” involves both localization and depth estimation. To isolate depth estimation, we explicitly provide the car’s bounding box. This approach enables fine-grained diagnostic insights into VFMs’ AVA-specific strengths and weaknesses. Detailed examples are illustrated in Figure 3, with further details specific below and in the Appendix.
@@ 3.2.1 Localization
Localization is a crucial foundational AVA, essential for various VQA questions. We curate 34.8K localization-focused image–question pairs sourced from diverse datasets: Objects365 shao2019objects365 and LVIS gupta2019lvis (open-domain), iNaturalist-2021 van2021benchmarking (birds and animals), and DIOR li2020object (remote-sensing). To ensure clarity, we include images with a single instance of the target object and exclude objects with bounding box area smaller than 0.2% of the image.
@@ 3.2.2 Counting
We collect 11.3K image-question pairs from five datasets: VQAv2 goyal2017making (open-domain), FSC ranjan2021learning (open-domain counting), CARPK hsieh2017drone (cars), Crowd Surveillance Dataset li2022video (people), and LVIS gupta2019lvis (open-domain). FSC, CARPK, and Crowd datasets explicitly cater to counting tasks, ensuring clear and distinct instance counts. LVIS instance masks are utilized to generate additional counting questions, complemented by counting-focused selections from VQAv2.
@@ 3.2.3 Spatial Reasoning
We curate 15.9K pairs from NYU-Depth V2 silberman2012indoor for indoor scenes with object segmentation, and LVIS gupta2019lvis and Objects365 shao2019objects365 for diverse open-domain scenarios. For each image, we select two non-overlapping objects—one marked with a blue bounding box (the reference object) and another with a red bounding box (the target). The model is then asked to determine the relative spatial position of the target object with respect to the reference one, choosing from multiple-choice options: top-left, top-right, bottom-left, and bottom-right. We ensure that the two boxes do not overlap either horizontally or vertically to avoid ambiguity.
@@ 3.2.4 Visual Attribute
Orientation. We compile 10.1K pairs from two specialized datasets, CURE-OR Temel2018_ICMLA and EgoOrientBench jung2024right , providing uncluttered images of objects from nine distinct orientations: front, back, left, right, top, front left, front right, back left, and back right.
Color. We curate 14K pair sourcing from from open-domain datasets, such as Objects365 shao2019objects365 and LVIS gupta2019lvis . To isolate the color recognition from object semantics or background clutter, we provide bounding boxes that tightly localize regions with minimal color variance, ensuring that the model focuses on RGB prediction.
Texture. We collect 13.2K pairs from four open-domain datasets: DTD cimpoi14describing , Kylberg kylberg2011kylberg , KTH-TIPS kth-tips_dataset and KTH-TIPS2 article with diverse texture types, such as striped, aluminum foil, and zigzagged.
Emotion. We gather 13.9K pairs from two publicly available datasets, RAF-DB li2019reliable and ExpW lian2020expression , covering 7 annotated human emotional states: surprise, neutral, disgust, fear, happy, sad, and angry.
@@ 3.2.5 Depth Estimation
We curate depth-focused image-question pairs from NYU-Depth V2 silberman2012indoor for indoor scenes and KITTI geiger2013vision for outdoor scenes.
@@ 3.2.6 Recognition
Action. We construct 15K pairs from the Moments in Time monfort2019moments , a video dataset with diverse human actions. We extract the middle frame from each 3-second video for 302 distinct actions
Fine-grained. We curate 9K pairs from three diverse datasets: CUB-200-2011 wah2011caltech (birds), iNaturalist-2021 van2021benchmarking (fungi, plants, animals), and FGVC-Aircraft maji2013fine (objects). We randomly select 50 species across fungi, plants, and animals from iNaturalist van2021benchmarking and 100 bird species from CUB, resulting in a total of 300 fine-grained classes.
Object. We curate 34.8K pairs from 4 datasets with diverse domains across 70 unique objects: Objects365 shao2019objects365 and LVIS gupta2019lvis (open-domain), iNaturalist-2021 van2021benchmarking (birds and animals), and DIOR li2020object (remote sensing). For each image, we provide a bounding box to eliminate the need for localization.
Scene. We curate 13.9K pairs from two datasets: Places434 zhou2017places (open-domain) and AID xia2017aid (remote sensing). The model is required to select the correct scene class from 30 randomly sampled options. The full set includes 464 classes spanning a wide range of scenes.
@@ 3.3 Quality Control and Dataset Statistics
To further improve the generalizability within each AVA, we aggregate data from multiple diverse datasets. This ensures that models encounter the same core visual ability across varied scenes, object types, and data distributions, making the assessment more robust and less dataset-specific. Lastly, image-question pairs are intentionally crafted or selected to be simple, clear, and explicitly focused on testing only one AVA at a time. In summary, AVA-Bench comprises 218K meticulously curated image-question pairs that robustly isolate individual AVAs, carefully control dataset balance and visibility, and systematically prevent annotation biases. Detailed statistics are provided in Figure 5, with additional details in Appendix.

### v3

@@ Abstract
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 1 Introduction
Figure 2: Visual Question Answering (VQA) often requires multiple atomic visual abilities to answer a question. When a model makes an incorrect prediction, it’s hard to determine whether it stems from a failure to capture all required AVAs or just a single critical one.
Vision Foundation Models (VFMs), pre-trained on large and diverse datasets, have become central to AI by providing transferable features for a wide range of downstream tasks [41, 2, 14, 5]. The variety of pre-training objectives and supervision signals has led to a proliferation of specialized VFMs—such as DINOv2 [70], CLIP [72], and SAM [44]—each excelling in distinct visual capabilities while often exhibiting interesting emergent properties [25, 6, 69]. Thus, establishing a systematic and effective evaluation protocol for VFMs becomes increasingly crucial.
Figure 3: AVA-Bench consists of 14 Atomic Visual Abilities that can be combined to address more complex visual reasoning tasks.
Existing evaluation protocols can generally be categorized into two groups. The first focuses on task-specific capabilities, typically attaching tailored heads to VFMs, followed by fine-tuning and evaluation on dedicated datasets such as ImageNet for classification [31, 59] and COCO for detection or segmentation [85, 3]. To better capture the diverse and complex perception challenges of the real world, recent studies advocate a more generic approach that leverages large language models (LLMs) as general-purpose heads, evaluating VFMs on broad Visual Question Answering (VQA) benchmarks [55, 117, 13].
While increasingly adopted, this generic protocol may suffer from two potential blind spots: (i) discrepancies between instruction tuning data and VQA test data lead to performance drops due to data mismatch rather than genuine visual limitations in VFMs, and (ii) existing VQA benchmarks typically require multiple visual abilities simultaneously, making it difficult to determine whether a failure arises from the absence of multiple abilities or merely a single critical one.
To this end, we introduce AVA-Bench, the first VFM evaluation benchmark explicitly designed to disentangle Atomic Visual Abilities (AVAs)—the fundamental visual capabilities that combine to solve complex visual reasoning tasks. For instance, answering typical VQA questions shown in Figure 2 necessitates integrating several AVAs.
@@
We systematically benchmark leading VFMs trained under diverse objectives and data (see LABEL:fig:_ava_eval), covering language-supervised (e.g., SigLIP-1/2 [88, 108], CLIP [72], InternVL-2.5 [10]), multimodal autoregressive (AIMv2 [21]), segmentation-supervised (SAM [44]), depth-supervised (MiDas [73]), contrastive self-supervised (DINOv2 [70]), and agglomerative models (RADIO [75]). We follow the standard protocol of adding an LLM on top, but fine-tune it separately for each AVA.
@@
  • We release a resource-efficient evaluation protocol with an open-source codebase to facilitate the development of the next generation of accountable and versatile VFMs.
Remark. We want to emphasize that AVA-Bench is designed to systematically evaluate VFMs (e.g., DINOv2) instead of MLLMs (e.g., LLaVA). MLLM in our study solely acts as a general-purpose prediction head for the underlying VFM, enabling a unified evaluation interface across tasks.
2 Vision Foundation Models and Evaluation
@@ 2.1 Vision Foundation Models (VFMs)
Figure 4: (a) Evaluation pipeline for AVA-Bench: The standard LLaVA-style two-stage training prepares the connector and LLM for VFM evaluation. For each AVA, only connector and LoRA is trained. (b) Overall statistics of AVA-Bench.
A wide range of architectures and learning strategies have been explored to develop vision foundation models (VFMs). In general, they fall into two categories based on the nature of their training data. On one hand, VFMs are pre-trained purely on visual data. For instance, the Vision Transformer (ViT)  [17] is trained on labeled images using the supervised signal to capture visual representations. DINOv2  [70], in contrast, adopts a self-supervised learning approach, enabling the model to learn those features without labeling. Some VFMs are designed for specific vision tasks—for example, SAM [44] specializes in open-vocabulary segmentation, while MiDaS [73] focuses on monocular depth estimation. RADIO  [75] instead introduces a multi-teacher distillation framework that unifies the strengths of different VFMs (e.g., CLIP, DINOv2, and SAM) into a single efficient student model.
Another family of VFMs leverages image-text pairs. CLIP  [72] uses contrastive loss to align image and text captions, while SigLIP  [88, 108] replaces the contrastive loss with a sigmoid one for efficient training. Unlike the conventional language-guided VFMs that process images and text separately, AIMv2  [21] integrates visual and textual understanding into a single auto-regressive framework, which is simple yet effective. Due to their different architectures and learning objectives, VFMs may have varying strengths and limitations in visual understanding.
@@ 2.2 LLM-based Evaluation for VFMs
Unlike the traditional paradigm, where different tasks (e.g. classification and segmentation) necessitate task-specific models [31, 2], the widespread adoption of Large Language Models (LLMs) as versatile interfaces has significantly shifted this paradigm [109, 97]. Reflecting this shift, recent studies advocate leveraging LLMs as general-purpose heads and evaluating VFMs on broad Visual Question Answering (VQA) benchmarks [86]. Specifically, following the LLaVA approach, these studies utilize a two-stage framework: (i) pre-training a connector between frozen LLM and VFMs with image-text pairs for feature alignment (ii) fine-tuning both the connector and the LLM using instruction-tuning data, while keeping VFMs frozen during both stages (see Figure 4 (a)). This LLM-based evaluation protocol has rapidly gained popularity as it closely mirrors the contemporary multimodal LLM setting and effectively captures diverse real-world perception challenges.
However, this evaluation protocol suffers from two potential blind spots. First, discrepancies may exist between instruction-tuning data and the test VQA datasets. Thus, a miss-prediction might arise from data mismatch rather than genuine visual deficiencies in VFMs. Second, typical VQA benchmarks typically require multiple visual abilities simultaneously to produce correct answers. This multi-ability requirement obscures the exact cause behind a model’s incorrect predictions. For instance, if a VFM is proficient in almost all visual abilities except orientation recognition (forward or backward), it would fail on the question in Figure 2 (left). In such cases, current evaluations only provide relative comparisons between VFMs without elucidating the specific missing capabilities. Thus, AVA-Bench aims to complement the LLM-based evaluation by pinpointing exactly where a VFM excels or falters, yielding a holistic understanding of their strengths and shortcomings.
@@ 3.1 Atomic Visual Abilities (AVAs)
To address the blind spots discussed earlier, we introduce AVA-Bench, the first systematic evaluation suite explicitly designed to disentangle 14 fundamental perceptual skills–Atomic Visual Abilities (AVAs)–for VFMs. AVAs are fundamental perceptual capabilities that can be combined to address more complex visual reasoning tasks. Rather than treating VQA as a monolithic task, we break it down into 14 fundamental perceptual abilities–such as counting, depth estimation, localization and spatial reasoning–each of which can be composed to answer more complex questions. For example, answering the question, “How many yellow dogs are facing backward on the left-hand side of the stop sign?”, requires multiple AVAs: counting, color recognition, localization, and spatial reasoning (see Figure 2). By explicitly defining the set of AVAs required to interpret a question, our benchmark quantitatively characterizes where a VFM excels or falters across each core ability.
The AVA selection is grounded in a thorough literature analysis, including compositional text-to-images benchmarks [37, 42, 99] and VQA questions [27, 1] (details in Appendix), focusing strictly on pure perceptual tasks and excluding non-perceptual reasoning skills (e.g., mathematical reasoning). Examples of each AVA can be found in Figure 3 and detailed definitions can be found in Appendix.
@@ 3.2 Dataset Curation
Constructing AVA-Bench required carefully isolating image–question pairs that specifically test individual AVAs. Existing MLLM and VQA datasets often blend multiple perceptual abilities, complicating direct assessments [87, 113, 107]. To achieve clear isolation, we assembled a comprehensive suite comprising image–question pairs from 26 diverse datasets, explicitly aligning each pair with a single targeted AVA. These datasets span a broad range of domains—including general scenes, wildlife (e.g., birds, fungi, plants), vehicles, indoor/outdoor settings, and remote-sensing imagery.
Given an AVA, image–question pairs in AVA-Bench were carefully designed or adapted to focus solely on that AVA. For instance, the question “What’s the depth of the car from the camera?” involves both localization and depth estimation. To isolate depth estimation, we explicitly provide the car’s bounding box. This approach enables fine-grained diagnostic insights into VFMs’ AVA-specific strengths and weaknesses. Examples are illustrated in Figure 3, with details below and in the Appendix.
@@ 3.2.1 Localization
We curate 34.8K localization-focused image–question pairs sourced from: Objects365 [78] and LVIS [30] (open-domain), iNat [91] (birds and animals), and DIOR [50] (remote-sensing). To ensure clarity, we include images with a single instance of the target object and exclude objects with small bounding boxes.
@@ 3.2.2 Counting
We collect 13.6K pairs from: VQAv2 [27] (open-domain), FSC [74] (open-domain counting), CARPK [34] (cars), Crowd Surveillance Dataset [49] (people), and LVIS [30] (open-domain). FSC, CARPK, and Crowd datasets explicitly cater to counting tasks, ensuring clear and distinct instance counts. LVIS instance masks are utilized to generate counting questions, complemented by counting-focused pairs from VQAv2.
@@ 3.2.3 Spatial Reasoning
We curate 11.5K pairs from NYU-Depth V2 [80] for indoor scenes, and LVIS [30] and Objects365 [78] for open-domain scenarios. For each image, we select two non-overlapping objects—one marked with a blue bounding box (the reference object) and another with a red bounding box (the target). The model is then asked to determine the relative spatial position of the target object with respect to the reference one, choosing from: top-left, top-right, bottom-left, and bottom-right.
@@ 3.2.4 Visual Attribute
Orientation. We compile 8.5K pairs from two specialized datasets, CURE-OR [83] and EgoOrientBench [40], providing uncluttered images of objects from nine distinct orientations: front, back, left, right, top, front left, front right, back left, and back right.
Color. We curate 14K pair sourcing from from open-domain datasets, such as Objects365 [78] and LVIS [30]. To isolate the color recognition from object semantics or background clutter, we provide bounding boxes that tightly localize regions with minimal color variance, ensuring that the model focuses on RGB prediction.
Texture. We curate 13.2K pairs from: DTD [15], Kylberg [46], KTH-TIPS and KTH-TIPS2 [65] with diverse texture types, such as striped, aluminum foil, and zigzagged.
Emotion. We gather 17.0K pairs fromRAF-DB [51] and ExpW [53], covering 7 annotated human emotional states: surprise, neutral, disgust, fear, happy, sad, and angry.
@@ 3.2.5 Depth Estimation
We curate depth-focused image-question pairs from NYU-Depth V2 [80] for indoor scenes and KITTI [24] for outdoor scenes.
@@ 3.2.6 Recognition
Action. We construct 15K pairs from Moments in Time [68], a video dataset with diverse human actions. We extract the middle frame from each video for 302 actions
Fine-grained. We curate 9K pairs from: CUB-200 [93] (birds), iNat-21 [91] (fungi, plants, animals), and Aircraft [64] (objects). We randomly select 50 species from iNat, 100 species from CUB and all aircraft classes, resulting in a total of 300 classes.
Object. We curate 44.9K pairs from 4 datasets with diverse domains across 70 unique objects: Objects365 [78] and LVIS [30] (open-domain), iNaturalist-2021 [91] (birds and animals), and DIOR [50] (remote sensing). For each image, we provide a bounding box to eliminate the need for localization.
Scene. We curate 13.9K pairs from two datasets: Places434 [116] (open-domain) and AID [101] (remote sensing). The model is required to select the correct scene class from 30 randomly sampled options. The full set includes 464 classes spanning a wide range of scenes.
[…]
@@ 3.3 Quality Control and Dataset Statistics
To improve the generalizability within each AVA, we aggregate data from diverse datasets. This ensures that models encounter the same core visual ability across varied scenes, object types, and data distributions, making the assessment more robust and less dataset-specific. Image-question pairs are intentionally crafted or selected to be simple, clear, and explicitly focused on testing only one AVA at a time. In summary, AVA-Bench comprises 218K meticulously curated image-question pairs that robustly isolate individual AVAs, carefully control dataset balance, visibility, and systematically prevent annotation biases. Statistics are provided in Figure 4 (b) , with more details in Appendix.